Metal cylindrical barrels are used in the disposal of hazardous wastes. They are 67 in. high and 32 in. across in diameter. Approximately how many cubic inches of waste can be stored in a barrel?
1
Expert's answer
2012-12-07T03:59:39-0500
Let's calculate the volume V of a barrel using the following formula:

V = π·R²·H,

where R and H are the radius and the height of a barrel respectfully. So,

V = π·(D/2)²·H = π·(32[inch]/2)²·67[inch] = 17152π [inch³] ≈ 53884.597 [inch³].

Therefore, there can be stored about 53884.597 cubic inches of waste.